OpenStudy (anonymous):

how do i simplify (8*10^7) (7*10^4)

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

8 , 7 and 10^7 , 10^4 are your like terms What do you think you should do with them ?

OpenStudy (doc.brown):

Everything is multiplied together in your question, so the order doesn't matter.\[a\times b\times c=c\times b\times a\]so \[(8\times 10^7)(7\times 10^4)=8\times 10^7\times 7\times 10^4\]@Shay17 was saying\[8\times 7\times 10^7\times 10^4\]

OpenStudy (anonymous):

Yes @doc.brown

OpenStudy (triciaal):

when you multiply numbers with the same base add the powers so you can simplify some more